# Beijing 100084, China

### v1

@@ Abstract
In this paper, we propose a novel function named Rate Distortion-in-Distortion (RDD) function as an extension of the classical rate-distortion (RD) function, where the expected distortion constraint is replaced by the Gromov-type distortion. This distortion, integral to the Gromov-Wasserstein (GW) distance, effectively defines the similarity in spaces of different dimensions without a direct metric between them. While our RDD function qualifies as an informational RD function, encoding theorems substantiate its status as an operational RD function, thereby underscoring its potential applicability in real-world source coding. Due to the high computational complexity associated with Gromov-type distortion, the RDD function cannot be solved analytically. Consequently, we develop an alternating mirror descent algorithm that significantly reduces computational complexity by employing decomposition, linearization, and relaxation techniques. Simulations on classical sources and different grids demonstrate the effectiveness of our algorithm. By examining the distinctions and connections between the RDD function and the RD function, we anticipate that RDD function will play a novel role in foreseeable future scenarios.
@@ I Introduction
The RDD function serves as an extension of the classical RD function. Like other RD extensions, it incorporates structural considerations into the rate-distortion problem, such as the perception term introduced by RDP function in the field of lossy compression [7]. Specifically, RDD adopts the Gromov-type distortion of two measurable spaces, rather than solely quantifying distances of the same dimension. This insight is provided by the GW distance, which offers a flexible and comprehensive framework for evaluating information processing systems, addressing limitations [8] in RD theory: the difficulty in defining a distortion measure across spaces of different dimensions, and the challenge of dealing with unknown distortion measures even within spaces of the same dimension.
@@
Our exploration of GW distance is motivated by the similarities between the Optimal Transport (OT) problem [9] and the RD problem. Specifically, the RD objective function partially corresponds to the entropy regularization in regularized OT, and the constraints on conditional probability in the RD function are related to the marginal distributions in OT problem [10]. As an important model in OT theory, GW distance provides a powerful method for evaluating discrepancies across distributions by measuring differences between distances rather than individual points [3]. Notably, the GW distance offers significant flexibility by enabling the comparison of probability distributions. In scenarios where the underlying metric spaces differ, GW distance captures structural and relational information that traditional distances may overlook [6]. Furthermore, when the distributions are supported on identical metric spaces, GW distance captures intrinsic geometric properties, thus offering valuable additional insights, such as shape analysis. Such functionality makes GW distance applicable in a wide range of mathematical and computational areas, including graph matching [6, 11, 12, 13], natural language processing [14], and machine learning [15, 16]. Although the RDD function, integrated with Gromov-type distortion, currently lacks specific application scenarios, we anticipate it would play a relevant role in the field of information theory.
@@
The RDD function is defined with mutual information as the objective function, positioning it naturally as an informational rate-distortion function. By invoking the coding theorem presented in [17], we can show that the RDD function corresponds to the minimum rate required for a specific coding task. Consequently, the RDD function holds operational significance for assessing the performance of real-world source coding.
@@
It is noteworthy that the RDD function involves quadratic constraints from the Gromov-type distortion, and thus constructing analytical solutions becomes vastly challenging. Therefore, an effective numerical algorithm is in high demand. In order to effectively reduce the quadratic arithmetic complexity, we develop an alternating mirror descent numerical algorithm, drawing inspiration from the computation of GW distance to decompose the Gromov-type distortion [18]. In addition, the RDD function involves an unknown marginal distribution, necessitating the relaxation of certain constraints and linearization of conditional probabilities to avoid complicated iterative operations caused by dual parameters in the Lagrangian function. By adapting the aforementioned techniques, we are able to reduce the arithmetic complexity from quartic to cubic.
@@ II Rate Distortion-in-Distortion Function
Let $(\mathcal{X},d_{\mathcal{X}},P_{\mathcal{X}})$ denote a metric measure space, and let $(\mathcal{Y},d_{\mathcal{Y}})$ represent a metric space, where $d_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}^{+}$ and $d_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ are two distances, while $P_{\mathcal{X}}$ is a probability measure on $\mathcal{X}$. We use $X$ and $Y$ to represent the random variables of spaces $\mathcal{X},\mathcal{Y}$, respectively. Clearly, the probability distribution of $X$ is $P_{\mathcal{X}}$, and for convenience, we will use $P_{X}$ to represent the distribution of $X$ throughout this paper. Given the unknown probability distribution on space $\mathcal{Y}$, the joint distribution of random variables $X$ and $Y$ refers to the product of the marginal distribution $P_{X}$ and the conditional probability $P_{Y|X}$, i.e., $P_{XY}=P_{X}\cdot P_{Y|X}$. We hereby introduce the RDD function as:
@@
where $q\geq 1$, and $I(X;Y)$ denotes the mutual information between $X$ and $Y$. The function $f$ measures the difference between the two metric $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$. Meanwhile, the joint distribution of $(X,Y,X^{\prime},Y^{\prime})$ is regarded as the product of two independent and identical distributions. Consequently, the expectation $\mathbb{E}$ in (1b) is taken with respect to the product of the joint distribution $P_{XY}\times P_{XY}$.
The commonly used element-wise loss function [13] $f$ includes Euclidean distance (MSE) $f(a,b)=(a-b)^{2}$ and KL-divergence $f(a,b)=a\log\frac{a}{b}-a+b$. In the rest of the paper, we adopt Euclidean distance to discuss theoretical properties and numerical algorithms due to its popularity in information theory. In such a case, the RDD function can be written as:
@@
$\iint_{(\mathcal{X}\times\mathcal{Y})^{2}}\!\!\!|d_{\mathcal{X}}^{q}(x,x^{\prime})-d_{\mathcal{Y}}^{q}(y,y^{\prime})|^{2}\mathop{}\!\mathrm{d}P_{XY}(x,y)\mathop{}\!\mathrm{d}P_{XY}(x^{\prime},y^{\prime}),$
Here we notice the similarity between the RDD function and the classical RD function. The classical RD function is obtained by minimizing the mutual information between the source and the reproduction subject to an average distortion constraint. Concretely, given a source random variable $X\in\mathcal{X}$ with probability distribution $P_{X}$, a reproduction $Y\in\mathcal{Y}$, the RD function is defined as [19, 1]:
@@
The distortion measure $d:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ here is primarily effective for spaces of identical dimensions. We refer to our function as rate distortion-in-distortion because we replace the distortion measure $d$ in RD function by the distortion-in-distortion $(d^{q}_{\mathcal{X}}-d^{q}_{\mathcal{Y}})^{2}$, which is inspired by the GW distance [3]:
$\mathrm{GW}^{2}_{q}((X,d_{\mathcal{X}}),(Y,d_{\mathcal{Y}}))=$
$\inf_{P_{XY}}\int_{(X\times Y)^{2}}\!\!\!\!\!\!\!\!\!\!\!\!(d^{q}_{\mathcal{X}}(x,x^{\prime})-d^{q}_{\mathcal{Y}}(y,y^{\prime}))^{2}\,\mathrm{d}P_{XY}(x,y)\,\mathrm{d}P_{XY}(x^{\prime},y^{\prime})$
The GW distance concerns the distance between two metric measure spaces, while the RD function focuses on compression. Leveraging insights from GW distance, the RDD function extends the application and definition of the RD problem, being able to handle spaces of different dimensions.
@@
See Appendix due to the space limitation. ∎
@@ III Coding Theorem
In this section, we use a coding theorem to show that $R_{G}(D)$ is the minimum number of bits required by a coding task for each source symbol. To start with, we define a code sequence as $\{(f_{n},\varphi_{n})\}_{n=1}^{\infty}$, where encoding function $f_{n}:\mathcal{X}^{n}\times\mathbb{R}\rightarrow\mathbb{Z}^{+}$ and decoding function $\varphi_{n}:\mathbb{Z}^{+}\times\mathbb{R}\rightarrow\mathcal{Y}^{n}$. Each $f_{n}$ represents an encoder, and each $\varphi_{n}$ represents a decoder. For two random vectors $(X_{1},\cdots,X_{n})$ and $(Y_{1},\cdots,Y_{n})$, they are said to satisfy the Gromov-type distortion constraint regarding distortion $D$, if
@@
satisfies for all positive integers $i\leq n$, where $P_{X_{i},Y_{i}}^{i}$ represents the joint distribution on the $i$-th variable pair $(X_{i},Y_{i})$. Now let $X_{1}$, $X_{2}$, $\cdots$, $X_{n}$ be independent and $P_{X}$-distributed random variables. Using the Theorem 3 in [17], we have the following coding theorem.
@@
For real numbers $R$ and $D$, $R\geq R_{G}(D)$ is a sufficient and necessary condition for the existence of a code sequence $\{(f_{n},\varphi_{n})\}_{n=1}^{\infty}$ and a sequence $\{U_{n}\}_{n=1}^{\infty}$ of random variables such that
@@
See Appendix due to the space limitation. ∎
The random variables $U_{1}$, $U_{2}$, $\cdots$, $U_{n}$ in the coding theorem serve as common randomness between the encoders and decoders. Since we can represent $f_{n}(X_{1},\cdots,X_{n},U_{n})$ using a variable-length code whose average length is approximately $H(f_{n}(X_{1},\cdots,X_{n},U_{n})|U_{n})$, we need at least $R_{G}(D)$ bits per source symbol to ensure that the Gromov-type distortion constraint regarding $D$ is satisfied.
The coding theorem presented in this section is more suitable for practical encoding design scenarios. Additionally, we provide an alternative perspective on the coding theorem in the Appendix.
@@ IV Alternating Mirror Descent Algorithm
Notably, the quadratic nature of the Gromov-type distortion constraint makes obtaining a closed-form solution for the RDD function theoretically challenging, even for Gaussian sources. To compute the RDD function in practical applications, effective algorithms are essential. However, the quartic arithmetic complexity from Gromov-type distortion presents significant computational challenges. To address these difficulties, we adapt the widely used Lagrangian multipliers method [10] with certain effective techniques to develop a novel numerical algorithm for the computation of the RDD function, which is referred to as Alternating Mirror Descent (AMD) algorithm.
In the sequel, we adopt a discretized formulation of the RDD function with discrete sources for computational convenience. Consider a memoryless source $X\in\mathcal{X}$ with reproduction $Y\in\mathcal{Y}$, where $\mathcal{X}\in\left\{x_{1},\cdots,x_{M}\right\},\mathcal{Y}\in\left\{y_{1},\cdots,y_{N}\right\}$ are finite and discrete alphabets. We denote $w_{ij}=P_{Y|X}(x_{i},y_{j})$ and $W$ is the matrix whose $(i,j)$ element is $w_{ij}$. Denote $D^{\mathcal{X}},D^{\mathcal{Y}}$ as the matrix whose $(i,j)$ element is $d_{\mathcal{X}}^{q}(x_{i},x^{\prime}_{j}),\ d_{\mathcal{Y}}^{q}(y_{i},y^{\prime}_{j})$ respectively. Let $p_{i}=P_{X}(x_{i}),r_{j}=P_{Y}(y_{j})$ and denote $E_{ij}:=D^{\mathcal{X}}_{ij}p_{j},C_{ij}:=E_{ij}p_{i}$ as the metrics for short. By introducing an auxiliary variable $\bm{r}$, the RDD function can be written in the discrete form as:
@@
IV-A Equivalent Transformation
Intuitively, the cost of evaluating (6c) is often unacceptable in the field of information science. Following the approach
in [12, 18], we can decompose the function $\mathcal{E}(D^{\mathcal{X}},D^{\mathcal{Y}},W)$ into one constant term $\widetilde{\mathcal{C}}_{1}$ and two quadratic terms of $W$ [^1] Unlike the GW distance, in the RDD function, one of the marginal distributions is unknown, leading to the decomposition of the Gromov-type distortion containing only a single constant term.:
@@ IV-A Equivalent Transformation
Moreover, similar to the treatment in [20], the constraints on marginal distributions in the second term of (6c), expressed as $\sum_{i=1}^{M}w_{ij}p_{i}=r_{j},\forall j$, can be relaxed during the subsequent Lagrangian analysis. This relaxation not only simplifies the formulation of the problem during the Lagrangian analysis, but also leads to a closed-form solution in the $r_{j}$ direction.
@@
See Appendix due to the space limitation. ∎

### v2

@@ Abstract
In this paper, we propose a novel function named Rate Distortion-in-Distortion (RDD) function as an extension of the classical rate-distortion (RD) function, where the expected distortion constraint is replaced by a Gromov-type distortion. This distortion, integral to the Gromov-Wasserstein (GW) distance, effectively defines the similarity in spaces of possibly different dimensions even without a direct metric between them. While the RDD function qualifies as an informational RD function, encoding theorems substantiate its status as an operational RD function, thereby underscoring its potential applicability in real-world source coding. Due to the high computational complexity associated with Gromov-type distortion, in general, the RDD function cannot be evaluated analytically. Consequently, we develop an alternating mirror descent algorithm that significantly reduces computational complexity by employing decomposition, linearization, and relaxation techniques. Numerical results on classical sources and different grids demonstrate the effectiveness of the developed algorithm. By exploring the relationship between the RDD function and the RD function, we suggest that the RDD function may have potential applications in future scenarios.
@@ I Introduction
The RDD function serves as an extension of the classical RD function. Like other RD extensions, it incorporates structural considerations into the rate-distortion problem, such as the perception term introduced by the RDP function in the field of lossy compression [7]. Specifically, RDD adopts the Gromov-type distortion that quantifies discrepancies between distance structures of two metric measure spaces, irrespective of their dimensions or explicit point-wise correspondence. This insight is provided by the GW distance, which offers a flexible and comprehensive framework for evaluating information processing systems, addressing limitations [8] in RD theory: the difficulty in defining a distortion measure across spaces of possibly different dimensions, and the challenge of dealing with unknown distortion measures even within spaces of the same dimension.
@@
Our exploration of the GW distance is motivated by similarities between the Optimal Transport (OT) problem [9] and the RD problem [10]. Specifically, the RD objective function corresponds to the entropy regularization in regularized OT, and the constraints on conditional probability in the RD function are related to the marginal distributions in OT problem [10]. As an important model in OT theory, the GW distance provides a powerful method for evaluating discrepancies across distributions by measuring differences between distances rather than individual points [3]. Notably, the GW distance offers significant flexibility by enabling the comparison of probability distributions. In scenarios where the underlying metric spaces differ, the GW distance captures structural and relational information that traditional distances may overlook [6]. Furthermore, when the distributions are supported on identical metric spaces, the GW distance captures intrinsic geometric properties, thus offering valuable additional insights. Such functionality makes the GW distance applicable in a wide range of mathematical and computational areas, including graph matching [6, 11, 12, 13], natural language processing [14], and machine learning [15, 16]. Although the RDD function, integrated with Gromov-type distortion, currently lacks specific application scenarios, we anticipate it will play a relevant role in the field of information theory.
@@
Building on these insights, we propose the RDD function with mutual information as its objective function, suggesting it may serve as a novel informational rate-distortion measure. By invoking the coding theorem presented in [17], we can show that the RDD function corresponds to the minimum rate required for a specific coding task. Consequently, the RDD function also holds operational utility for assessing the performance of real-world source coding.
@@
It is noteworthy that the RDD function involves quadratic constraints from the Gromov-type distortion, and thus constructing its analytical solutions becomes vastly challenging. Therefore, an effective numerical algorithm is in demand. In order to effectively reduce the quartic arithmetic complexity, we develop an alternating mirror descent numerical algorithm, drawing inspiration from the computation of the GW distance to decompose the Gromov-type distortion [18]. In addition, the RDD function involves an unspecified marginal distribution, necessitating the relaxation of certain constraints and linearization of conditional probabilities to avoid complicated iterative operations caused by dual parameters in the Lagrangian function. By adapting the aforementioned techniques, we are able to reduce the arithmetic complexity from quartic to cubic.
@@
For readers unfamiliar with GW distance, we provide a minimal primer: for metric spaces $(\mathcal{X},d_{\mathcal{X}})$ and $(\mathcal{Y},d_{\mathcal{Y}})$ with measures $P_{\mathcal{X}}$ and $P_{\mathcal{Y}}$, GW distance is formulated as [3]:
$\mathrm{GW}^{2}_{q}((\mathcal{X},d_{\mathcal{X}}),(\mathcal{Y},d_{\mathcal{Y}}))=$
$\inf_{P_{\mathcal{X}\mathcal{Y}}}\iint_{(\mathcal{X}\times\mathcal{Y})^{2}}\!\!\!\!\!\!\!\!\!\!\!\!|d^{q}_{\mathcal{X}}(x,x^{\prime})-d^{q}_{\mathcal{Y}}(y,y^{\prime})|^{2}\,\mathrm{d}P_{\mathcal{X}\mathcal{Y}}(x,y)\,\mathrm{d}P_{\mathcal{X}\mathcal{Y}}(x^{\prime},y^{\prime}),$
where $P_{\mathcal{X}\mathcal{Y}}$ couples $P_{\mathcal{X}}$ and $P_{\mathcal{Y}}$. Subsequently, we will present the Gromov-type distortion metric as given in (3).
@@ II Rate Distortion-in-Distortion Function
Let $(\mathcal{X},d_{\mathcal{X}},P_{\mathcal{X}})$ denote a metric measure space, and let $(\mathcal{Y},d_{\mathcal{Y}})$ represent a metric space, where $d_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}^{+}$ and $d_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ are two distances, while $P_{\mathcal{X}}$ is a probability measure on $\mathcal{X}$. We use $X$ and $Y$ to represent the random variables in spaces $\mathcal{X},\mathcal{Y}$, respectively. Clearly, the probability distribution of $X$ is $P_{\mathcal{X}}$, and for convenience, we will use $P_{X}$ to represent the distribution of $X$ throughout this paper. The joint distribution of random variables $X$ and $Y$ refers to the product of the marginal distribution $P_{X}$ and the conditional probability $P_{Y|X}$, i.e., $P_{XY}=P_{X}\cdot P_{Y|X}$. We hereby introduce the RDD function as:
@@
where $q\geq 1$, and $I(X;Y)$ denotes the mutual information between $X$ and $Y$. The function $f$ measures the difference between the two metrics $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$. Meanwhile, the joint distribution of $(X,Y,X^{\prime},Y^{\prime})$ is regarded as the product of two independent and identical distributions. Consequently, the expectation $\mathbb{E}$ in (1b) is taken with respect to the product of the joint distribution $P_{XY}\times P_{XY}$.
Commonly used element-wise loss function [13] $f$ includes Euclidean distance (MSE) $f(a,b)=(a-b)^{2}$ and KL-divergence $f(a,b)=a\log\frac{a}{b}-a+b$. In the rest of the paper, we adopt Euclidean distance to study theoretical properties and numerical algorithms due to its popularity in practice. In such a case, the RDD function can be written as:
@@
$\iint_{(\mathcal{X}\times\mathcal{Y})^{2}}\!|d_{\mathcal{X}}^{q}(x,x^{\prime})-d_{\mathcal{Y}}^{q}(y,y^{\prime})|^{2}\mathop{}\!\mathrm{d}P_{XY}(x,y)\mathop{}\!\mathrm{d}P_{XY}(x^{\prime},y^{\prime}).$ (3)
Here we notice the similarity between the RDD function and the classical RD function. The classical RD function is obtained by minimizing the mutual information between the source and the reproduction subject to an average distortion constraint. Concretely, given a source $X\in\mathcal{X}$ with probability distribution $P_{X}$, and a reproduction $Y\in\mathcal{Y}$, the RD function is defined as [19, 1]:
@@
The distortion measure $d:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ here is primarily used to measure the discrepancy between $X$ and $Y$. We refer to our function as rate distortion-in-distortion, as it substitutes the distortion measure $d(x,y)$ in the RD function with a structural distortion $(d^{q}_{\mathcal{X}}-d^{q}_{\mathcal{Y}})^{2}$, thereby enforcing consistency between the intrinsic geometries of $\mathcal{X}$ and $\mathcal{Y}$. Leveraging insights from the GW distance, the RDD function extends the application and definition of the RD problem, being able to handle different spaces of possibly different dimensions.
@@
See Appendix due to space limitation. ∎
@@ III Coding Theorem
In this section, we use a coding theorem to show that $R_{G}(D)$ is the minimum number of bits required by a coding task for each source symbol. To start with, we define a code sequence as $\{(f_{n},\varphi_{n})\}_{n=1}^{\infty}$, where $f_{n}:\mathcal{X}^{n}\times\mathbb{R}\rightarrow\mathbb{Z}^{+}$ and $\varphi_{n}:\mathbb{Z}^{+}\times\mathbb{R}\rightarrow\mathcal{Y}^{n}$. Each $f_{n}$ represents an encoder, and each $\varphi_{n}$ represents a decoder. For two random vectors $(X_{1},\cdots,X_{n})$ and $(Y_{1},\cdots,Y_{n})$, they are said to satisfy the Gromov-type distortion constraint regarding distortion $D$, if
@@
is satisfied for all positive integers $i\leq n$, where $P_{X_{i},Y_{i}}^{i}$ represents the joint distribution on the $i$-th pair $(X_{i},Y_{i})$. Now let $X_{1}$, $X_{2}$, $\cdots$, $X_{n}$ be i.i.d. $P_{X}$-distributed random variables. Using Theorem 3 in [17], we have the following coding theorem.
@@
For $R$ and $D$, $R\geq R_{G}(D)$ is a sufficient and necessary condition for the existence of a code sequence $\{(f_{n},\varphi_{n})\}_{n=1}^{\infty}$ and a sequence $\{U_{n}\}_{n=1}^{\infty}$ of random variables such that
@@
See Appendix due to space limitation. ∎
The random variables $U_{1}$, $U_{2}$, $\cdots$, $U_{n}$ in Theorem 1 serve as common randomness between encoders and decoders. Since we can represent $f_{n}(X_{1},\cdots,X_{n},U_{n})$ using a variable-length code whose average length is approximately $H(f_{n}(X_{1},\cdots,X_{n},U_{n})|U_{n})$, we need at least $R_{G}(D)$ bits per source symbol to ensure that the Gromov-type distortion constraint regarding $D$ is satisfied.
The coding theorem presented herein is applicable to the scenario where a single stream of source symbols is to be coded. On the other hand, one may also adopt a somewhat conceived model with two independent source streams, and establish a coding theorem for the RDD function, as provided in the Appendix.
@@ IV Alternating Mirror Descent Algorithm
Notably, the quadratic nature of the Gromov-type distortion constraint makes obtaining a closed-form solution for the RDD function theoretically challenging, even for Gaussian sources. To compute the RDD function in applications, effective algorithms are essential. We adopt the widely used Lagrangian multipliers method [10] with certain effective techniques to develop a novel numerical algorithm for the computation of the RDD function, which is referred to as Alternating Mirror Descent (AMD) algorithm.
In the sequel, we consider a discretized formulation of the RDD function with discrete sources for computational convenience. Consider a source $X\in\mathcal{X}$ with reproduction $Y\in\mathcal{Y}$, where $\mathcal{X}\in\left\{x_{1},\cdots,x_{M}\right\},\mathcal{Y}\in\left\{y_{1},\cdots,y_{N}\right\}$ are finite discrete alphabets. We denote $w_{ij}=P_{Y|X}(x_{i},y_{j})$ and $W$ is the matrix whose $(i,j)$ element is $w_{ij}$. Denote $D^{\mathcal{X}},D^{\mathcal{Y}}$ as matrices whose $(i,j)$ element is $d_{\mathcal{X}}^{q}(x_{i},x^{\prime}_{j}),\ d_{\mathcal{Y}}^{q}(y_{i},y^{\prime}_{j})$ respectively. Let $p_{i}=P_{X}(x_{i}),r_{j}=P_{Y}(y_{j})$ and denote $E_{ij}:=D^{\mathcal{X}}_{ij}p_{j},C_{ij}:=E_{ij}p_{i}$ as the metrics for short. By introducing an auxiliary variable $\bm{r}$, the RDD function can be written in discrete form as:
@@
IV-A Equivalent Relaxation
Directly computing the Gromov-type distortion in (7c) incurs quartic complexity, which is computationally prohibitive. Following the approach in [12, 18], we can decompose $\mathcal{E}(D^{\mathcal{X}},D^{\mathcal{Y}},W)$ into one constant term $\widetilde{\mathcal{C}}_{1}$ and two quadratic terms of $W$ [^1] Unlike the GW distance, in the RDD function, one of the marginal distributions is unknown, leading to the decomposition of the Gromov-type distortion containing only a single constant term.:
@@ IV-A Equivalent Relaxation
Moreover, similar to the treatment in [20], the constraints on marginal distributions in the second term of (7c), expressed as $\sum_{i=1}^{M}w_{ij}p_{i}=r_{j},\forall j$, can be relaxed during the subsequent Lagrangian analysis. This relaxation simplifies the formulation of the problem during the Lagrangian analysis, and further leads to a closed-form solution in the $r_{j}$ direction.
@@
See Appendix due to space limitation. ∎